OpenStudy (anonymous):

Partial Derivative A manufacturer estimates that the annual output at a certain factory is given by Q(K,L)=(42K^0.8)(L^0.2) units, where K is the capital expenditure in thousands of dollars and L is the size of the labor force in worker-hours. Suppose that the current capital expenditure is $240000 and the labor level is 560 worker-hours.

OpenStudy (tkhunny):

Where is the partial derivative? So far, just substitute the given values.

OpenStudy (anonymous):

I tried that but it did not work. And our homework is over partial derivatives so I was unsure what we were supposed to do lol

OpenStudy (anonymous):

Find the current marginal productivity of capital QK

OpenStudy (tkhunny):

Aha!! A question. That we can do. Have you considered the Total Derivative?

OpenStudy (anonymous):

Sorry I forgot that part haha... No I have not, should I try it? (33.6K^-.2)(.2L^-.8) is the derivative, right?

OpenStudy (anonymous):

I tried it and it didn't work

OpenStudy (tkhunny):

In one variable. \(A(t) = \pi r^{2}\) \(\dfrac{dA}{dr} = 2\pi r\) or the differential form \(dA = 2\pi r\;dr\) This is all we are doing. We are just doing it in multiple dimensions. A partial deriviative re Q(K,L)=(42K^0.8)(L^0.2) \(\dfrac{\partial Q}{\partial K} = 48\cdot 0.8\cdot K^{0.8 - 1.0}\cdot L^{0.2}\) or the differential form. \(\partial Q = \left(48\cdot 0.8\cdot K^{0.8 - 1.0}\cdot L^{0.2}\right)\;\partial K\) Do that for L and add them up.

OpenStudy (anonymous):

So do I just plug in 240,000 for K and 560 for L?

OpenStudy (tkhunny):

1) "Plug In" doesn't actually mean anything. 2) No. Just above, you were asked to find \(\dfrac{dQ}{dL}\). Did you do that?

OpenStudy (anonymous):

No...

OpenStudy (tkhunny):

Well...

OpenStudy (anonymous):

I'm not sure what I'm supposed to do lol

OpenStudy (tkhunny):

You are finding a partial derivative with respect to L. Treat K as a constant and find the 1st Derivative.

OpenStudy (anonymous):

2(-.16(K)^(-1.2))(L^.2)?

OpenStudy (tkhunny):

Why did the K exponent change? That's no good. Treat K like a constant. Just carry it along for the ride. Why did you add 1 to the L exponent. This is a derivative. It should go DOWN by 1.

OpenStudy (anonymous):

(.8K^-.2)(.2L^-.8)2K

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

(42K^0.8)(L^0.2) Change Nothing: (42K^0.8) Find the Derivative: (L^0.2)